# On Krylov Complexity as a Probe of the Quantum Mpemba Effect

### v1

@@ 1 Introduction
Recently, Krylov complexity [^1] For comprehensive reviews, see [38, 39]. has emerged as a powerful and physically transparent diagnostic of quantum dynamics and chaos [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. By quantifying how rapidly a quantum state or operator explores its accessible Hilbert space, Krylov complexity encodes both the geometric and dynamical aspects of relaxation that are often inaccessible to conventional observables. Within this framework, recent work [40] on a certain spin-chain system with global $U(1)$ symmetry reported clear Mpemba-like behavior in the symmetric (diagonal) component of Krylov complexity, obtained by neglecting contributions from the asymmetric (off-diagonal) sectors. Interestingly, inclusion of the full asymmetric contribution tends to smooth out these features, suggesting that coherence between symmetry sectors plays a subtle yet nontrivial role in shaping relaxation hierarchies. This observation points toward a deeper connection between complexity growth, symmetry constraints, and the emergence of anomalous relaxation in quantum many-body systems.
@@ 2 Krylov Complexity and the Quantum Mpemba Effect without global symmetry
In this section, we explore how Krylov complexity [^2] Throughout this paper, we examine the Krylov complexity of states, following [63, 64, 65]. can serve as a probe of the quantum Mpemba effect (QME) in systems without global symmetries. Following [37], we study the mixed-field Ising chain with the Hamiltonian
@@
Figure 1: Early-time growth of Krylov complexity for different tilt angles $\theta$. Strongly tilted states initially grow more slowly but eventually overtake weakly tilted ones, a signature of Mpemba-like behavior.
@@ 3 Krylov Complexity and the Quantum Mpemba Effect in $U(1)$-Symmetric Systems
This possibility was explicitly demonstrated in Ref. [40] for the Aubry–André model [71], where the total complexity showed only a qualitative late-time separation, while the symmetric component exhibited robust Mpemba-like crossings, making it a more sensitive diagnostic of the QME. In particular, a modified definition of the symmetric complexity was introduced,
@@
To examine whether this behavior persists in interacting systems, we now turn to other $U(1)$-symmetric spin chains. We first consider the next-nearest-neighbor XXZ Hamiltonian [34, 37],
which is chaotic and possesses a global $U(1)$ symmetry generated by the total magnetization along the $z$-axis,
$Q=\sum_{i=1}^{L}\sigma_{i}^{z}.$ (3.12)
To analyze the dynamics, we compute the total and symmetric Krylov complexities for TFS with different tilt parameters. To highlight the crossings, we study the difference

### v2

@@ 1 Introduction
Recently, Krylov complexity [^1] For comprehensive reviews, see [38, 39]. has emerged as a powerful and physically transparent diagnostic of quantum dynamics and chaos [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64]. By quantifying how rapidly a quantum state or operator explores its accessible Hilbert space, Krylov complexity encodes both the geometric and dynamical aspects of relaxation that are often inaccessible to conventional observables. Within this framework, recent work [40] on a certain spin-chain system with global $U(1)$ symmetry reported clear Mpemba-like behavior in the symmetric (diagonal) component of Krylov complexity, obtained by neglecting contributions from the asymmetric (off-diagonal) sectors. Interestingly, inclusion of the full asymmetric contribution tends to smooth out these features, suggesting that coherence between symmetry sectors plays a subtle yet nontrivial role in shaping relaxation hierarchies. This observation points toward a deeper connection between complexity growth, symmetry constraints, and the emergence of anomalous relaxation in quantum many-body systems.
@@ 2 Krylov Complexity and the Quantum Mpemba Effect without global symmetry
In this section, we explore how Krylov complexity [^2] Throughout this paper, we examine the Krylov complexity of states, following [65, 66, 67, 68]. can serve as a probe of the quantum Mpemba effect (QME) in systems without global symmetries. Following [37], we study the mixed-field Ising chain with the Hamiltonian
@@
We also examine the next-nearest-neighbor XXZ Hamiltonian [34, 37]
which is known to be chaotic and possesses a global $U(1)$ symmetry generated by the total magnetization along the $z$-axis,
$Q=\sum_{i=1}^{L}\sigma_{i}^{z}.$ (2.3)
In addition to the continuous $U(1)$ symmetry, this model also exhibits two discrete symmetries: parity and a global $Z_{2}$ symmetry. The role of these symmetries in shaping Krylov complexity will be explored in more detail in the next section. Here, however, we focus on the fully symmetry-broken case in order to isolate the intrinsic dynamical features of complexity growth. To this end, we set $J_{1}=J_{2}$ and consider the following Hamiltonian:
$H=\sum_{i=1}^{L-1}\big(\sigma_{i}^{x}\sigma_{i+1}^{x}+\Delta_{y}\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta_{z}\sigma_{i}^{z}\sigma_{i+1}^{z}\big)+\sum_{i=1}^{L-2}\big(\sigma_{i}^{x}\sigma_{i+2}^{x}+\Delta_{y}\sigma_{i}^{y}\sigma_{i+2}^{y}+\Delta_{z}\sigma_{i}^{z}\sigma_{i+2}^{z}\big)+h\sum_{i=2}^{L-1}\sigma_{i}^{z}+h^{\prime}(\sigma_{1}^{z}-\sigma_{L}^{z}).$ (2.4)
Here, $\Delta_{y}\neq 1$ explicitly breaks the global $U(1)$ symmetry, while the boundary and linear field terms break the remaining discrete symmetries, ensuring that the system is fully non-symmetric.
@@
Figure 1: Early-time growth of Krylov complexity for different tilt angles $\theta$ in next-nearest-neighbor XXZ Hamiltonian (Right) and mixed-field Ising chain (Left). Strongly tilted states initially grow more slowly but eventually overtake weakly tilted ones, a signature of Mpemba-like behavior.
@@ 3 Krylov Complexity and the Quantum Mpemba Effect in $U(1)$-Symmetric Systems
This possibility was explicitly demonstrated in Ref. [40] (see also [75, 76]) for the Aubry–André model [74], where the total complexity showed only a qualitative late-time separation, while the symmetric component exhibited robust Mpemba-like crossings, making it a more sensitive diagnostic of the QME. In particular, a modified definition of the symmetric complexity was introduced,
@@
To examine whether this behavior persists in interacting systems, we now turn to other $U(1)$-symmetric spin chains. We first consider the next-nearest-neighbor XXZ Hamiltonian (2.2). To analyze the dynamics, we compute the total and symmetric Krylov complexities for TFS with different tilt parameters. To highlight the crossings, we study the difference